Janny opened a savings account and deposited $300.00 as principal. The account earns 6% interest, compounded annually. What is the balance after 6 years? *

1
Expert's answer
2021-12-19T15:44:14-0500

"A=P(1+r/100)^n=300(1+6/100)^6=300\times 1.06^6=\$425.56"